OpenStudy (anonymous):

solve the following equation by factorisation method 20x^3=5x

OpenStudy (anonymous):

20x^3=5x subtract 5x from both sides 20x^3-5x=0 factor out 5x 5x(4x^2-1)=0 so 5x=0 x=0 and 4x^2-1=0 4x^2=1 x^2=1/4 \[x=\sqrt{1/4}\] x=1/2, -1/2

OpenStudy (anonymous):

thank you!